the table gives the probbability distribution of the number of dvds sold per day at a music store. what is the probability of 5 or more dvds being sold on any given day?
0-4: .1875
5-9:?
10-14:.4375
15-20:.125

Answer by stanbon(75887)

the table gives the probability distribution of the number of dvds sold per day at a music store. what is the probability of 5 or more dvds being sold on any given day?
0-4: .1875
5-9:?
10-14:.4375
15-20:.125
------
? = 1 - (0.1875+0.4375+0.1250) = 0.2500
------
Ans: 1 - 0.1875 = 0.8125
---------------
Cheers,
Stan H.
-------------